OpenStudy (anonymous):

if y=u^4 + 5u^2, where u= x^5+2x^2+1, find dy/dx. Leave your answer in terms of u and x.

OpenStudy (adi3):

@AaronAndyson

OpenStudy (adi3):

@ganeshie8

OpenStudy (rational):

use chain rule

OpenStudy (anonymous):

I got this: 4(x^5+2x^2+1)^3(5x^4+4x) + 5(2(x^5+2x^2+1)(5x^4+4x)

OpenStudy (rational):

thats right! but the question clearly tells us to leave the answer in terms of "u" and "x"

OpenStudy (rational):

y=u^4 + 5u^2 dy/dx = 4u^3 du/dx + 5(2u) du/dx = (4u^3+10u) du/dx = (4u^3+10u)(5x^4+4x)

OpenStudy (anonymous):

oohh.. so we sub it

OpenStudy (rational):

leave it like that. i don't think you need to replace "u"

OpenStudy (anonymous):

thanks!!! :)